Cheetahs can accelerate to a speed of 19.8 m/s in 2.65 s and can continue to accelerate to reach a top speed of 28.8 m/s. Assume the acceleration is constant until the top speed is reached and is zero thereafter. Let the +x direction point in the direction the cheetah runs. Express the cheetahs top speed Vop in miles per hour (mi/h) mi/h Starting from a crouched position, how much time tae does it take a cheetah to reach its top speed and what distance d does it travel in that time? accel d= If a cheetah sees a rabbit 128.0 m awav. how much time fom.l will t take to reach the rabbit, assuming the rabbit does not
uop- mi/h Starting from a crouched position, how much time faccel does it take a cheetah to reach its top speed and what distance d does it travel in that time? laccel = 2 If a cheetah sees a rabbit 128.0 m away, how much time fotal will it take to reach the rabbit, assuming the rabbit does not move and the cheetah starts from rest? total
Answer #1

1 m/s =2.23694 mi/h

a)

Vtop=28.8*2.23694=64.424 mi/h

b)

From

V=Vo+at

since it starts from rest Vo=0

=>19.8=0+a(2.65)

a=7.4717 m/s2

From

=>28.8 = 0+(7.4717)taccel

taccel=3.8545 s

From

d=Vot+(1/2)at2

d=0+(1/2)*7.4717*3.85452

d=55.5 m

c)

Distance traveled after reaching constant speed

D=128-55.5 =72.5 m

time taken to travel 72.5 m

tcons = D/V =72.5/28.8=2.517 s

Total time taken to travel 128 m is

ttotal=3.8545+2.517

ttotal=6.372 s